# Deep Learning, Machine Learning - Digital Signal and Image Processing: From Theory to Application

### v1

@@ 9.2.1 Types of Noise
Several types of noise commonly affect digital images, each with unique statistical properties. Understanding the characteristics of each noise type is crucial for selecting the appropriate restoration method [16].
@@
Gaussian noise is one of the most common types of noise and is often introduced by electronic components during image acquisition. It is modeled by a normal distribution with a mean $\mu$ and variance $\sigma^{2}$ [118]. The probability density function (PDF) of Gaussian noise is given by:
@@
Salt-and-pepper noise, also known as impulse noise, is characterized by randomly occurring white and black pixels in the image. This noise is caused by sudden disturbances during image acquisition, such as transmission errors or malfunctioning camera sensors [26].
@@
Poisson noise, also known as shot noise, is caused by the random nature of photon counting during image acquisition, especially in low-light environments. This noise follows a Poisson distribution, and its variance is proportional to the intensity of the signal. Unlike Gaussian noise, Poisson noise has a more significant impact on regions with higher intensity [118].
@@ 9.4 Image Restoration Techniques
Image restoration involves recovering an image that has been degraded by a known or unknown process. The goal of restoration is to reverse the degradation as much as possible, thereby improving the image quality. Various techniques are used in image restoration, each with its strengths and limitations depending on the type of degradation and noise present [11]. In this section, we will explore inverse filtering, Wiener filtering, and regularization methods, providing detailed explanations and examples to help beginners grasp these concepts thoroughly.
@@ 9.4.1 Inverse Filtering
Inverse filtering is one of the simplest methods for image restoration, and it involves reversing the effect of a known degradation process [16]. This technique assumes that the degradation model is known and that the degradation can be modeled as a linear process in the frequency domain.
@@ 9.4.2 Wiener Filtering
Wiener filtering provides a more robust solution for image restoration when noise is present. Unlike inverse filtering, Wiener filtering takes into account the statistical properties of both the noise and the original image, providing an optimal solution in terms of minimizing the mean square error [164, 153].
@@ 9.4.3 Regularization Methods
Regularization methods are particularly useful when the degradation model is incomplete or when there is significant noise in the image. These techniques add a constraint or penalty to the restoration process to prevent overfitting to the noise. One common regularization method is Tikhonov regularization, which introduces a smoothness constraint on the restored image [17].
@@
Each of the image restoration techniques discussed has its advantages and limitations [11]:
@@ 9.5 Blind Restoration Techniques
Blind deconvolution is a technique used to recover a sharp image from a blurred one, without knowing the exact nature of the blur [219]. The process involves estimating both the original image and the blurring function simultaneously. This is a challenging problem because it is underdetermined—there are many possible combinations of original images and degradation functions that could produce the same blurred image [101].
@@
Applications: Blind deconvolution is used in various fields such as medical imaging, astronomical imaging, and photography, where the degradation function is unknown or difficult to measure [101].
@@ 9.6 Image Interpolation and Reconstruction
Image interpolation is a key technique used when enlarging images or reconstructing missing pixel values. Interpolation methods estimate new pixel values by using surrounding pixels. This process is crucial in applications such as image scaling, rotation, and image reconstruction from sparse data [194]. In this section, we explore three common interpolation techniques: nearest-neighbor interpolation, bilinear interpolation, and bicubic interpolation.
@@ 9.6.1 Nearest-Neighbor Interpolation
Nearest-neighbor interpolation is the simplest form of interpolation. When enlarging an image, this method assigns the value of the nearest pixel in the original image to each pixel in the enlarged image. While this approach is computationally efficient, it often produces blocky or pixelated images, as it does not smooth out transitions between neighboring pixel values [173].
@@
Applications: Nearest-neighbor interpolation is often used in applications where simplicity and speed are more important than visual quality, such as enlarging low-resolution images or generating thumbnails [173].
@@ 9.6.2 Bilinear Interpolation
Bilinear interpolation is a more advanced technique compared to nearest-neighbor interpolation [185]. It smooths the transitions between pixels by averaging the values of the four closest pixels. This method results in smoother images and is commonly used when a higher-quality enlargement is required.
@@
Applications: Bilinear interpolation is widely used in image resizing, where maintaining a smoother appearance is essential, such as in digital zooming, rotation, and texture mapping [185].
@@ 9.6.3 Bicubic Interpolation
Bicubic interpolation is an even more advanced technique that considers the 16 nearest pixels to compute the new pixel value, using cubic polynomials to calculate the weights. Bicubic interpolation produces higher-quality results than both nearest-neighbor and bilinear interpolation, especially for large scaling factors [49].
@@
Image segmentation and edge detection are foundational techniques in digital image processing, used for partitioning an image into meaningful regions or identifying boundaries between different objects [58]. These methods are critical for applications such as object detection, image recognition, medical imaging, and computer vision. By breaking an image into distinct regions, segmentation simplifies the task of analyzing specific parts of the image, enabling more effective object detection and pattern recognition [74, 14, 145].
@@ 10.1 Basic Concepts of Image Segmentation
Image segmentation is the process of dividing an image into multiple segments or regions, each of which represents a specific part of the image. The goal of segmentation is to simplify the image into something more meaningful and easier to analyze by isolating objects or regions of interest [32].
@@
  • Object Detection: Identifying objects such as people, cars, or animals in images or videos [74].
  • Medical Imaging: Segmenting different tissue types in medical scans (e.g., MRI or CT scans) to aid in diagnosis [145].
  • Scene Understanding: Partitioning different areas in a landscape image, such as the sky, mountains, and trees [104].
@@
  • Thresholding-based Segmentation: This method involves partitioning the image based on pixel intensity values. Thresholding is widely used because of its simplicity and efficiency [131].
@@
  • Region-based Segmentation: This method divides the image into regions based on similarities in color, intensity, or texture [83].
  • Clustering-based Segmentation: Techniques like $k$-means clustering or mean shift group pixels into clusters based on feature similarity [176].
@@ 10.2.1 Global Thresholding
In global thresholding, a single threshold value is applied to the entire image. The goal is to choose a threshold $T$ that effectively separates the objects from the background [100]. Mathematically, this can be expressed as:
@@ 10.2.2 Local Thresholding
Local thresholding (also known as adaptive thresholding) divides the image into smaller regions and applies a different threshold to each region. This method is particularly useful when the lighting conditions vary across the image, as it can account for local variations in intensity [81].
@@
  • The adaptiveThreshold() function applies local thresholding. The threshold is calculated based on the mean intensity of the neighboring pixels in an $11\times 1$ 11 block. The constant $C=2$ is subtracted from the mean to fine-tune the threshold.
@@ 10.2.3 Otsu’s Method for Automatic Thresholding
Otsu’s method is a widely used global thresholding technique that automatically determines the optimal threshold value by minimizing the intra-class variance, which is the variance within the foreground and background regions. It works by maximizing the variance between these two classes [10].
Otsu’s method computes a histogram of the image and then evaluates every possible threshold to find the one that maximizes the separation between the two classes. The advantage of Otsu’s method is that it does not require manual selection of a threshold value, making it ideal for automating the segmentation process [110].
@@
    For each possible threshold $T$, divide the pixels into two classes: foreground (pixels $\geq T$) and background (pixels $<T$).
@@ 10.4 Edge Detection Techniques
Edge detection is a fundamental technique in image processing that aims to identify significant local changes in intensity, which typically correspond to object boundaries, texture changes, or other important features in an image. Effective edge detection is crucial for applications such as image segmentation, object recognition, and computer vision. In this section, we will explore three important edge detection techniques: the Sobel operator, the Prewitt operator, and the Canny edge detection algorithm. Each method is introduced with clear explanations, practical examples, and a discussion of its advantages and limitations [223, 27].
@@ 10.4.1 Sobel Operator
The Sobel operator is a simple and widely-used method for detecting edges in images by computing the gradient magnitude. It works by applying a pair of convolution kernels to the image-one to detect changes in the horizontal direction ($G_{x}$) and one to detect changes in the vertical direction ($G_{y}$). The gradient magnitude at each pixel is then calculated from these two components [27].
@@ 10.4.2 Prewitt Operator
The Prewitt operator is similar to the Sobel operator in that it detects edges by calculating the gradient magnitude in an image. However, the Prewitt operator uses slightly different convolution kernels that provide a simpler approximation of the derivative [27].
@@ 10.4.3 Canny Edge Detection
The Canny edge detection algorithm is one of the most advanced and widely used edge detection techniques. It provides robust edge detection by addressing many of the limitations of simpler gradient-based methods like the Sobel and Prewitt operators. The Canny algorithm uses a multi-stage process that includes noise reduction, gradient calculation, non-maximum suppression, and hysteresis thresholding [213].
@@ 10.4.4 Comparison of Edge Detection Techniques
Each edge detection technique discussed has its strengths and weaknesses, making them suitable for different applications [223]:
@@ 10.5 Region-based Segmentation
Region-based segmentation is a fundamental technique in image processing, where the goal is to divide an image into regions that correspond to different objects or features. This type of segmentation is based on the similarity of pixel properties, such as intensity, color, or texture. Unlike edge-based segmentation, which relies on detecting discontinuities between regions, region-based methods group pixels into regions based on homogeneity [83]. In this section, we will explore two common techniques: region growing and split-and-merge segmentation.
@@ 10.5.1 Region Growing
Region growing is a straightforward technique in region-based segmentation. The idea is to start with one or more seed pixels and grow a region by adding neighboring pixels that have similar properties. The growth process continues until no more pixels can be added to the region based on the similarity criteria [1].
@@
  • Medical imaging (e.g., tumor detection, organ segmentation) [194].
  • Remote sensing and satellite imagery analysis [154].
  • Object detection in computer vision [193].
@@ 10.5.2 Split-and-Merge Segmentation
Split-and-merge segmentation is a more structured approach that divides the image into regions based on homogeneity. This method recursively splits the image into smaller regions and merges regions that meet a consistency criterion. The technique is often implemented using a quadtree structure, where each node represents a rectangular region of the image [31].
@@
In split-and-merge segmentation, the image can be represented as a quadtree, where each node of the tree represents a region of the image. The root node represents the entire image, and each node has four children corresponding to the four quadrants [210].
@@
  • Satellite image analysis and land cover classification [154].
  • Object recognition in high-resolution images [194].
@@ 10.6 Watershed Algorithm
The watershed algorithm is a popular technique in image segmentation that is based on the topography of the image. The algorithm interprets the grayscale intensity of an image as a surface or landscape, with pixel intensities corresponding to elevations. The basic idea is to treat the image as a topographic surface and ”flood” it with water, where the basins correspond to segments [91].
@@
Watershed $\text{Watershed boundaries}=\text{argmax}\left(G(x,y)\right)$
@@
  • Medical image segmentation (e.g., separating cells or tissues) [145].
  • Object detection and separation in cluttered scenes [193].
  • Satellite image analysis and terrain mapping [154].
@@
Image compression is an essential aspect of digital image processing, allowing us to reduce the storage and transmission requirements of images while maintaining an acceptable level of visual quality. The primary goal of compression is to represent an image with fewer bits by eliminating redundancies, without significantly compromising the quality of the image [58]. There are two main types of compression: lossless and lossy. In lossless compression, the original image can be perfectly reconstructed from the compressed data, whereas lossy compression achieves higher compression ratios by allowing some loss of image data, which may lead to a degradation in visual quality.
@@ 11.1 Basic Principles of Image Compression
Image compression works by exploiting three primary types of redundancy [44, 201]:
@@ 11.2 Lossless Compression Techniques
Lossless compression ensures that the original image can be perfectly reconstructed from the compressed image [216]. This is crucial in applications where any loss of information is unacceptable, such as medical imaging, technical drawings, and archival photography [88, 168, 89]. Two popular lossless compression techniques are Huffman coding and run-length encoding (RLE) [128, 33].

### v2

@@ 9.2.1 Types of Noise
Several types of noise commonly affect digital images, each with unique statistical properties. Understanding the characteristics of each noise type is crucial for selecting the appropriate restoration method [boncelet2009image].
@@
Gaussian noise is one of the most common types of noise and is often introduced by electronic components during image acquisition. It is modeled by a normal distribution with a mean $\mu$ and variance $\sigma^{2}$ [luisier2010image]. The probability density function (PDF) of Gaussian noise is given by:
@@
Salt-and-pepper noise, also known as impulse noise, is characterized by randomly occurring white and black pixels in the image. This noise is caused by sudden disturbances during image acquisition, such as transmission errors or malfunctioning camera sensors [chan2005salt].
@@
Poisson noise, also known as shot noise, is caused by the random nature of photon counting during image acquisition, especially in low-light environments. This noise follows a Poisson distribution, and its variance is proportional to the intensity of the signal. Unlike Gaussian noise, Poisson noise has a more significant impact on regions with higher intensity [luisier2010image].
@@ 9.4 Image Restoration Techniques
Image restoration involves recovering an image that has been degraded by a known or unknown process. The goal of restoration is to reverse the degradation as much as possible, thereby improving the image quality. Various techniques are used in image restoration, each with its strengths and limitations depending on the type of degradation and noise present [banham1997digital]. In this section, we will explore inverse filtering, Wiener filtering, and regularization methods, providing detailed explanations and examples to help beginners grasp these concepts thoroughly.
@@ 9.4.1 Inverse Filtering
Inverse filtering is one of the simplest methods for image restoration, and it involves reversing the effect of a known degradation process [boncelet2009image]. This technique assumes that the degradation model is known and that the degradation can be modeled as a linear process in the frequency domain.
@@ 9.4.2 Wiener Filtering
Wiener filtering provides a more robust solution for image restoration when noise is present. Unlike inverse filtering, Wiener filtering takes into account the statistical properties of both the noise and the original image, providing an optimal solution in terms of minimizing the mean square error [robinson1967principles, pratt1972generalized].
@@ 9.4.3 Regularization Methods
Regularization methods are particularly useful when the degradation model is incomplete or when there is significant noise in the image. These techniques add a constraint or penalty to the restoration process to prevent overfitting to the noise. One common regularization method is Tikhonov regularization, which introduces a smoothness constraint on the restored image [bouhamidi2007sylvester].
@@
Each of the image restoration techniques discussed has its advantages and limitations [banham1997digital]:
@@ 9.5 Blind Restoration Techniques
Blind deconvolution is a technique used to recover a sharp image from a blurred one, without knowing the exact nature of the blur [you1999blind]. The process involves estimating both the original image and the blurring function simultaneously. This is a challenging problem because it is underdetermined—there are many possible combinations of original images and degradation functions that could produce the same blurred image [levin2009understanding].
@@
Applications: Blind deconvolution is used in various fields such as medical imaging, astronomical imaging, and photography, where the degradation function is unknown or difficult to measure [levin2009understanding].
@@ 9.6 Image Interpolation and Reconstruction
Image interpolation is a key technique used when enlarging images or reconstructing missing pixel values. Interpolation methods estimate new pixel values by using surrounding pixels. This process is crucial in applications such as image scaling, rotation, and image reconstruction from sparse data [thevenaz2000image]. In this section, we explore three common interpolation techniques: nearest-neighbor interpolation, bilinear interpolation, and bicubic interpolation.
@@ 9.6.1 Nearest-Neighbor Interpolation
Nearest-neighbor interpolation is the simplest form of interpolation. When enlarging an image, this method assigns the value of the nearest pixel in the original image to each pixel in the enlarged image. While this approach is computationally efficient, it often produces blocky or pixelated images, as it does not smooth out transitions between neighboring pixel values [rukundo2012nearest].
@@
Applications: Nearest-neighbor interpolation is often used in applications where simplicity and speed are more important than visual quality, such as enlarging low-resolution images or generating thumbnails [rukundo2012nearest].
@@ 9.6.2 Bilinear Interpolation
Bilinear interpolation is a more advanced technique compared to nearest-neighbor interpolation [smith1981bilinear]. It smooths the transitions between pixels by averaging the values of the four closest pixels. This method results in smoother images and is commonly used when a higher-quality enlargement is required.
@@
Applications: Bilinear interpolation is widely used in image resizing, where maintaining a smoother appearance is essential, such as in digital zooming, rotation, and texture mapping [smith1981bilinear].
@@ 9.6.3 Bicubic Interpolation
Bicubic interpolation is an even more advanced technique that considers the 16 nearest pixels to compute the new pixel value, using cubic polynomials to calculate the weights. Bicubic interpolation produces higher-quality results than both nearest-neighbor and bilinear interpolation, especially for large scaling factors [fadnavis2014images].
@@
Image segmentation and edge detection are foundational techniques in digital image processing, used for partitioning an image into meaningful regions or identifying boundaries between different objects [gonzalez2002digital]. These methods are critical for applications such as object detection, image recognition, medical imaging, and computer vision. By breaking an image into distinct regions, segmentation simplifies the task of analyzing specific parts of the image, enabling more effective object detection and pattern recognition [hoeser2020object, bhanu2000adaptive, patil2013medical].
@@ 10.1 Basic Concepts of Image Segmentation
Image segmentation is the process of dividing an image into multiple segments or regions, each of which represents a specific part of the image. The goal of segmentation is to simplify the image into something more meaningful and easier to analyze by isolating objects or regions of interest [cheng2001color].
@@
  • Object Detection: Identifying objects such as people, cars, or animals in images or videos [hoeser2020object].
  • Medical Imaging: Segmenting different tissue types in medical scans (e.g., MRI or CT scans) to aid in diagnosis [patil2013medical].
  • Scene Understanding: Partitioning different areas in a landscape image, such as the sky, mountains, and trees [li2009towards].
@@
  • Thresholding-based Segmentation: This method involves partitioning the image based on pixel intensity values. Thresholding is widely used because of its simplicity and efficiency [mustaqeem2012efficient].
@@
  • Region-based Segmentation: This method divides the image into regions based on similarities in color, intensity, or texture [karthick2014survey].
  • Clustering-based Segmentation: Techniques like $k$-means clustering or mean shift group pixels into clusters based on feature similarity [saxena2017review].
@@ 10.2.1 Global Thresholding
In global thresholding, a single threshold value is applied to the entire image. The goal is to choose a threshold $T$ that effectively separates the objects from the background [lee1990comparative]. Mathematically, this can be expressed as:
@@ 10.2.2 Local Thresholding
Local thresholding (also known as adaptive thresholding) divides the image into smaller regions and applies a different threshold to each region. This method is particularly useful when the lighting conditions vary across the image, as it can account for local variations in intensity [jiang2003adaptive].
@@
  • The adaptiveThreshold() function applies local thresholding. The threshold is calculated based on the mean intensity of the neighboring pixels in an $11\times 11$ block. The constant $C=2$ is subtracted from the mean to fine-tune the threshold.
@@ 10.2.3 Otsu’s Method for Automatic Thresholding
Otsu’s method is a widely used global thresholding technique that automatically determines the optimal threshold value by minimizing the intra-class variance, which is the variance within the foreground and background regions. It works by maximizing the variance between these two classes [bangare2015reviewing].
Otsu’s method computes a histogram of the image and then evaluates every possible threshold to find the one that maximizes the separation between the two classes. The advantage of Otsu’s method is that it does not require manual selection of a threshold value, making it ideal for automating the segmentation process [liu2009otsu].
@@
    For each possible threshold $T$, divide the pixels into two classes: foreground (pixels $\geq T$) and background (pixels <T<T).
@@ 10.4 Edge Detection Techniques
Edge detection is a fundamental technique in image processing that aims to identify significant local changes in intensity, which typically correspond to object boundaries, texture changes, or other important features in an image. Effective edge detection is crucial for applications such as image segmentation, object recognition, and computer vision. In this section, we will explore three important edge detection techniques: the Sobel operator, the Prewitt operator, and the Canny edge detection algorithm. Each method is introduced with clear explanations, practical examples, and a discussion of its advantages and limitations [ziou1998edge, chaple2015comparisions].
@@ 10.4.1 Sobel Operator
The Sobel operator is a simple and widely-used method for detecting edges in images by computing the gradient magnitude. It works by applying a pair of convolution kernels to the image-one to detect changes in the horizontal direction ($G_{x}$) and one to detect changes in the vertical direction ($G_{y}$). The gradient magnitude at each pixel is then calculated from these two components [chaple2015comparisions].
@@ 10.4.2 Prewitt Operator
The Prewitt operator is similar to the Sobel operator in that it detects edges by calculating the gradient magnitude in an image. However, the Prewitt operator uses slightly different convolution kernels that provide a simpler approximation of the derivative [chaple2015comparisions].
@@ 10.4.3 Canny Edge Detection
The Canny edge detection algorithm is one of the most advanced and widely used edge detection techniques. It provides robust edge detection by addressing many of the limitations of simpler gradient-based methods like the Sobel and Prewitt operators. The Canny algorithm uses a multi-stage process that includes noise reduction, gradient calculation, non-maximum suppression, and hysteresis thresholding [xu2017canny].
@@ 10.4.4 Comparison of Edge Detection Techniques
Each edge detection technique discussed has its strengths and weaknesses, making them suitable for different applications [ziou1998edge]:
@@ 10.5 Region-based Segmentation
Region-based segmentation is a fundamental technique in image processing, where the goal is to divide an image into regions that correspond to different objects or features. This type of segmentation is based on the similarity of pixel properties, such as intensity, color, or texture. Unlike edge-based segmentation, which relies on detecting discontinuities between regions, region-based methods group pixels into regions based on homogeneity [karthick2014survey]. In this section, we will explore two common techniques: region growing and split-and-merge segmentation.
@@ 10.5.1 Region Growing
Region growing is a straightforward technique in region-based segmentation. The idea is to start with one or more seed pixels and grow a region by adding neighboring pixels that have similar properties. The growth process continues until no more pixels can be added to the region based on the similarity criteria [adams1994seeded].
@@
  • Medical imaging (e.g., tumor detection, organ segmentation) [thevenaz2000image].
  • Remote sensing and satellite imagery analysis [pritt2017satellite].
  • Object detection in computer vision [terven2023comprehensive].
@@ 10.5.2 Split-and-Merge Segmentation
Split-and-merge segmentation is a more structured approach that divides the image into regions based on homogeneity. This method recursively splits the image into smaller regions and merges regions that meet a consistency criterion. The technique is often implemented using a quadtree structure, where each node represents a rectangular region of the image [chen1979segmentation].
@@
In split-and-merge segmentation, the image can be represented as a quadtree, where each node of the tree represents a region of the image. The root node represents the entire image, and each node has four children corresponding to the four quadrants [wu1993adaptive].
@@
  • Satellite image analysis and land cover classification [pritt2017satellite].
  • Object recognition in high-resolution images [thevenaz2000image].
@@ 10.6 Watershed Algorithm
The watershed algorithm is a popular technique in image segmentation that is based on the topography of the image. The algorithm interprets the grayscale intensity of an image as a surface or landscape, with pixel intensities corresponding to elevations. The basic idea is to treat the image as a topographic surface and ”flood” it with water, where the basins correspond to segments [kornilov2018overview].
@@
$\text{Watershed boundaries}=\text{argmax}\left(G(x,y)\right)$
@@
  • Medical image segmentation (e.g., separating cells or tissues) [patil2013medical].
  • Object detection and separation in cluttered scenes [terven2023comprehensive].
  • Satellite image analysis and terrain mapping [pritt2017satellite].
@@
Image compression is an essential aspect of digital image processing, allowing us to reduce the storage and transmission requirements of images while maintaining an acceptable level of visual quality. The primary goal of compression is to represent an image with fewer bits by eliminating redundancies, without significantly compromising the quality of the image [gonzalez2002digital]. There are two main types of compression: lossless and lossy. In lossless compression, the original image can be perfectly reconstructed from the compressed data, whereas lossy compression achieves higher compression ratios by allowing some loss of image data, which may lead to a degradation in visual quality.
@@ 11.1 Basic Principles of Image Compression
Image compression works by exploiting three primary types of redundancy [dhawan2011review, vijayvargiya2013survey]:
@@ 11.2 Lossless Compression Techniques
Lossless compression ensures that the original image can be perfectly reconstructed from the compressed image [yang2005overview]. This is crucial in applications where any loss of information is unacceptable, such as medical imaging, technical drawings, and archival photography [kivijarvi1998comparison, rojatkar2015image, koff2006overview]. Two popular lossless compression techniques are Huffman coding and run-length encoding (RLE) [moffat2019huffman, chetan2017performance].